# Shuffle Products, Degenerate Affine Hecke Algebras, and Quantum Toda Lattice

### v1

@@ Appendix A Technical lemmas
The following result about symmetric polynomials is used in Proposition 5.7. Define $p^{\hbar}_{a}(\vec{\Omega})\in\mathbf{C}[\hbar][\Omega_{1},\ldots,\Omega_{k}]^{S_{k}}$ by
@@
where we understand each term in the product as in Remark 4.8.
@@
Therefore, the differential of $\varphi_{\hbar}$ at zero is invertible. By the proof of [GanGinzburg, Lemma 2.1], it is an isomorphism. In particular, the polynomials $\{p^{\hbar}_{a}(\vec{\Omega})\}$ generate the ring of symmetric functions. ∎
In what follows, we use the notations of Section 4 for $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes(S^{\bullet}\mathbf{C}^{N}\otimes S^{\bullet}(\mathbf{C}^{N})^{*})$ (more precisely, for its Mickelsson algebra).
@@
extending the one (1.6) of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})$. By Remark 4.3 we have $\varpi_{N}(P)=P$, therefore,
@@
The proof of the following proposition is a word-for-word repetition of [MolevTransvectorAlgebras, Theorem 3.1] adapted to the setting of this paper.
@@
Denote by $\alpha=N+1$. Recall the trick of Example 3.6. Denote by $T_{1\ldots N}^{1\ldots j\mapsto\alpha\ldots N}(u)$ the quantum minor with values in $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N+1})$ corresponding to the indices
@@
in the notations of (1.9), where $\alpha$ is on the $j$-th place. By Proposition 1.1 and (1.12), we have
@@
Therefore, we can study the image of the left-hand side in the quotient $\mathfrak{n}_{-}\backslash\mathrm{U}_{\hbar}(\mathfrak{gl}_{N+1})$, where $\mathfrak{n}_{-}\subset\mathfrak{gl}_{N}$. Recall the definition of the quantum minor (1.9) and the extremal projector notation from Section 4. Using the properties (4.6) and pushing the terms of $\mathfrak{n}_{-}$ to the left, we obtain
@@
$\bar{\phi}_{a}v_{a}P=\sum_{i=1}^{a}(w_{a}-w_{1})\ldots(w_{a}-w_{i-1})(w_{a}-w_{i+1}+\hbar)\ldots(w_{a}-w_{a-1}+\hbar)P\phi_{a}v_{a}P.$
@@ Appendix B Dual representation
In particular, the analog of the Cayley-Hamilton theorem of Proposition 3.7 reads as
@@
which follows directly from [MolevYangians, Theorem 7.2.1]. Similarly to Proposition 1.3, one can show that
@@
Recall the notations for the parabolic restriction of Section 4. As in (4.10), we can define
@@
where $s=0,1\ldots$ and $\{j_{1},\ldots,j_{r}\}$ is the complementary subset to $\{i_{1},\ldots,i_{s}\}$ in $\{1\ldots,i-1\}$. By the trick of Example 3.6 and [MolevTransvectorAlgebras, (3.3)], it is essentially given by the action of the extremal projector
@@
as in 4.6.
Recall the positive part ${}^{N}\tilde{\mathcal{A}}_{\hbar}^{+}$ of the algebra of difference operators ${}^{N}\tilde{\mathcal{A}}_{\hbar}$ from Section 4.4. We have an analog of Theorem 4.11
@@
Next, recall the Kostant-Whittaker reduction of Section 5, in particular the Kirillov projector $P_{\psi}$. The following equality can be proved analogously to [KalmykovYangians, Proposition 5.20].
@@
The following is an analog of Proposition 5.5.
@@
is an isomorphism. Therefore, it will be enough to relate the action of $\Omega^{*}$ to that of the Kirillov projector. By pushing the $\mathfrak{gl}_{N}$-terms to the left or to the right and using of Theorem 5.2 as in [KalmykovYangians, Proposition 6.18], we deduce
@@
The following is an analog of Proposition 5.7 and 5.8.
@@
Recall the definition of the non-positive part $\mathrm{Y}_{-N,N}(\mathfrak{gl}_{2})^{\leq 0}$ of the shifted Yangian from 2.13. The following is an analog of Theorem 5.9.
@@
Finally, recall the Miura transform of 6.4:
@@
Using (B.6), the following theorem can be proved analogously to Theorem 6.7.
@@
The map (B.8) coincides with the GKLO map from Theorem 2.16.
The following result is used in Section 7.2.
@@
where $e^{-}_{k}$ is the antisymmetrizer (2.6) acting on $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes((\mathbf{C}^{N})^{*})^{\otimes k}$, hence on its Kostant-Whittaker reduction. The case $k=1$ is obvious, and for the step, it follows from the proof of Proposition B.1 that
@@
By $\mathfrak{H}^{-\hbar}_{k}$-equivariance of Proposition B.1 and Proposition 2.7, the right-hand side is equal to the action of the element
@@
Finally, consider the topological setting of Section 7.1. Let
@@
[…]
@@
for $i=-1,0$. The left (resp. the right) action of $H^{\bullet}_{G}(\mathrm{pt})$ on $H^{\bullet}_{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}(\mathrm{Gr}^{\omega_{1}})$ is via the Chern classes of $\mathcal{V}_{0}$ (resp. $\mathcal{V}_{-1}$).
@@
The following is an analog of Proposition 7.13. Recall the isomorphism of Proposition B.1. Also, since the loop rotation is trivial on $\mathcal{Q}$, we have
@@
The rest of the proof goes as in Proposition 7.13 (recall that the right module structure on the target is given by (B.2)). ∎
[…]

### v2

@@ Appendix A Technical lemmas
The following result is used in the proof of Proposition˜2.3. Recall that
$G^{\theta}(z,w)=\frac{\theta(z+w)\theta^{\prime}(0)}{\theta(z)\theta(w)},$
where $\theta(z)$ is any skew-symmetric function satisfying the three-term relation (2.1).
Lemma A.1.
We have
$\prod_{i=1}^{l}G^{\theta}(u-z_{i},\kappa)=\sum_{i=1}^{l}G^{\theta}(u-z_{i},l\kappa)\cdot\prod_{\begin{subarray}{c}j=1\ldots l\\ j\neq i\end{subarray}}G^{\theta}(z_{i}-z_{j},\kappa).$
Proof.
We can show it by induction on $l$. The base $l=1$ is tautological. For the step, let us multiply both sides by $G^{\theta}(u-z_{l+1},\kappa)$. The three-term relation and skew-symmetricity of $\theta(t)$ imply
$\frac{\theta(u-z_{i}+l\kappa)\theta(u-z_{l+1}+\hbar)}{\theta(u-z_{i})\theta(l\kappa)\theta(u-z_{l+1})\theta(\kappa)}=$
$=$ $\frac{\theta(u-z_{l+1}+l\kappa+\kappa)\theta(z_{l+1}-z_{i}+l\kappa)}{\theta(u-z_{l+1})\theta(l\kappa+\kappa)\theta(z_{l+1}-z_{i})\theta(l\kappa)}+\frac{\theta(z_{i}-z_{l+1}+\kappa)\theta(u-z_{i}+l\kappa+\kappa)}{\theta(z_{i}-z_{l+1})\theta(\kappa)\theta(u-z_{i})\theta(l\kappa+\kappa)}$
by mutiplying by common denominator and taking
$x=\frac{l\kappa+\kappa+2u-z_{i}-z_{l+1}}{2},\qquad y=\frac{l\kappa-\kappa-z_{i}+z_{l+1}}{2},$
$z=\frac{l\kappa+\kappa-z_{i}+z_{l+1}}{2},\qquad w=\frac{l\kappa+\kappa+z_{i}-z_{l+1}}{2}$
in the notations of (2.1). Then the statement for $l+1$ follows from the induction assumption for $u=z_{l+1}$. ∎
The following result about symmetric polynomials is used in Proposition˜5.7. Define $p^{\hbar}_{a}(\vec{\Omega})\in\mathbf{C}[\hbar][\Omega_{1},\ldots,\Omega_{k}]^{S_{k}}$ by
@@
where we understand each term in the product as in Remark˜4.8.
@@
Therefore, the differential of $\varphi_{\hbar}$ at zero is invertible. By the proof of [GG02, Lemma 2.1], it is an isomorphism. In particular, the polynomials $\{p^{\hbar}_{a}(\vec{\Omega})\}$ generate the ring of symmetric functions. ∎
In what follows, we use the notations of Section˜4 for $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes(S^{\bullet}\mathbf{C}^{N}\otimes S^{\bullet}(\mathbf{C}^{N})^{*})$ (more precisely, for its Mickelsson algebra).
@@
extending the one (1.6) of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})$. By Remark˜4.3 we have $\varpi_{N}(P)=P$, therefore,
@@
The proof of the following proposition is a word-for-word repetition of [Mol02, Theorem 3.1] adapted to the setting of this paper.
@@
Denote by $\alpha=N+1$. Recall the trick of Example˜3.6. Denote by $T_{1\ldots N}^{1\ldots j\mapsto\alpha\ldots N}(u)$ the quantum minor with values in $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N+1})$ corresponding to the indices
@@
in the notations of (1.9), where $\alpha$ is on the $j$-th place. By Proposition˜1.1 and (1.12), we have
@@
[…]
@@
$\bar{\phi}_{a}v_{a}P=\sum_{i=1}^{a}\prod_{k=1}^{i-1}(w_{a}-w_{k})\cdot\prod_{k=i+1}^{a}(w_{a}-w_{k}+\hbar)\cdot P\phi_{i}v_{i}P.$
@@ Appendix B Dual representation
In particular, the analog of the Cayley-Hamilton theorem of Proposition˜3.7 reads as
@@
which follows directly from [Mol07, Theorem 7.2.1]. Similarly to Proposition˜1.3, one can show that
@@
Recall the notations for the parabolic restriction of Section˜4. As in (4.10), we can define
@@
where $s=0,1\ldots$ and $\{j_{1},\ldots,j_{r}\}$ is the complementary subset to $\{i_{1},\ldots,i_{s}\}$ in $\{1\ldots,i-1\}$. By the trick of Example˜3.6 and [Mol02, (3.3)], it is essentially given by the action of the extremal projector
@@
as in ˜4.6.
Recall the positive part $^{N}\tilde{\mathcal{A}}_{\hbar}^{+}$ of the algebra of difference operators $^{N}\tilde{\mathcal{A}}_{\hbar}$ from Section˜4.4. We have an analog of Theorem˜4.11
@@
Next, recall the Kostant-Whittaker reduction of Section˜5, in particular the Kirillov projector $P_{\psi}$. The following equality can be proved analogously to [Kal25, Proposition 5.20].
@@
The following is an analog of Proposition˜5.5.
@@
is an isomorphism. Therefore, it will be enough to relate the action of $\Omega^{*}$ to that of the Kirillov projector. By pushing the $\mathfrak{gl}_{N}$-terms to the left or to the right and using of Theorem˜5.2 as in [Kal25, Proposition 6.18], we deduce
@@
The following is an analog of Proposition˜5.7 and ˜5.8.
@@
Recall the definition of the non-positive part $\mathrm{Y}_{-N,N}(\mathfrak{gl}_{2})^{\leq 0}$ of the shifted Yangian from ˜2.13. The following is an analog of Theorem˜5.9.
@@
Finally, recall the Miura transform of ˜6.4:
@@
Using (B.6), the following theorem can be proved analogously to Theorem˜6.7.
@@
The map (B.8) coincides with the GKLO map from Theorem˜2.16.
The following result is used in Section˜7.2.
@@
where $e^{-}_{k}$ is the antisymmetrizer (2.6) acting on $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes((\mathbf{C}^{N})^{*})^{\otimes k}$, hence on its Kostant-Whittaker reduction. The case $k=1$ is obvious, and for the step, it follows from the proof of Proposition˜B.1 that
@@
By $\mathfrak{H}^{-\hbar}_{k}$-equivariance of Proposition˜B.1 and Proposition˜2.7, the right-hand side is equal to the action of the element
@@
Finally, consider the topological setting of Section˜7.1. Let
@@
(B.9) ${\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{R})\otimes(\mathbf{C}^{N})^{*}/\!\!/_{\psi}\,N_{+}^{R}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tilde{\varpi}_{N}}$${N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N^{L}_{-}\,\backslash\!\!\backslash_{-\psi^{*}}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi^{*}}N_{-}^{R}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{R})\otimes\mathbf{C}^{N}/\!\!/_{\psi^{*}}\,N_{-}^{R}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\sim}$${H^{\bullet}_{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}(\mathrm{Gr}^{\omega_{1}},\underline{\mathbf{C}}_{\mathrm{Gr}^{\omega_{1}}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}_{\bullet}(\mathrm{Gr})}$
@@
for i=1,0i=-1,0. The left (resp. the right) action of $H^{\bullet}_{G}(\mathrm{pt})$ on $H^{\bullet}_{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}(\mathrm{Gr}^{\omega_{1}})$ is via the Chern classes of $\mathcal{V}_{0}$ (resp. $\mathcal{V}_{-1}$).
@@
The following is an analog of Proposition˜7.13. Recall the isomorphism of Proposition˜B.1. Also, since the loop rotation is trivial on $\mathcal{Q}$, we have
@@
The rest of the proof goes as in Proposition˜7.13 (recall that the right module structure on the target is given by (B.2)). ∎